# Deciphering interventional dynamical causality from non-intervention systems

### v1

@@ Results
We validated that IEE was a promising indicator for assessing the transmission dynamics of infectious diseases. Our study collected the daily confirmed COVID-19 cases from all $47$ prefectures in Japan, spanning a duration of 1209 days from January 16, 2020 to May 8, 2023 (fig. S11).
@@
First, we calculated the IEE indices for IntDC among different prefectures. IEE effectively ranked the influence across prefectures, with the top five affected areas by Tokyo being Kanagawa, Chiba, Saitama, Aichi, and Osaka. Remarkably, this ranking coincides with the geographical proximity and socio-economic connections to Tokyo. Kanagawa, Chiba and Saitama are in the same metropolitan area with Tokyo, while Aichi and Osaka are far from Tokyo but connected to Tokyo by expressway and Shinkansen.
@@
Then, IEE was validated to coincide with the effective distance matrix $D^{\text{COVID}}$ between prefectures. The $D^{\text{COVID}}\in\mathbb{R}^{47\times 47}$ was non-symmetric and designed by incorporating factors such as geometric distances, human mobility, population sizes, and infectious rates across prefectures (based on the gravity model  (?, ?), see Supplementary Text and fig. S11 for details). The element $D_{ij}^{\text{COVID}}$ served as a benchmark for quantifying COVID-19 transmission from prefecture $i$ to $j$. The IEE values demonstrated a strong linear correlation with $\ln D^{\text{COVID}}$, particularly when assessing the causality from Tokyo/Osaka to other prefectures (Figs. 6(A-B)). This correlation remained significant for other regions such as Aichi, Hokkaido, Fukuoka, and Okinawa (see fig. S12).
@@
Further, IEE outperformed ObsDC indices in practicality. Notably, the absolute value of the Pearson correlation coefficient (PCC) between IEE and $\ln D^{\text{COVID}}$ was $0.724$ substantially higher than those obtained by GC $0.161$, TE $0.284$, and CCM $0.568$ (average over specific prefecture, see the second to last column in Table 3 and the boxplot in Fig. 6(C)). Table 3 presented more detailed comparisons to support the effectiveness of IEE (with Tokyo for specific illustration in fig. S13).
@@
Figure 6: Application of the IEE on the COVID-19 transmission in Japan and the gene regulatory network (GRN) of circadian rhythms. (A) shows that IEE values from Tokyo to other prefectures are highly linear-correlated with $\ln D^{\text{COVID}}$, where $D^{\text{COVID}}$ is the effective distance obtained from the gravity model in consideration of various factors. (B) is the result for IEE values from Osaka to the other prefectures. (C) is the boxplot for $|\text{PCC}(c_{ij}^{\text{Inf}},\ln D_{ij}^{\text{COVID}})|$ with average over fixed $i$, where PCC is the Pearson correlation coefficient and $c^{\text{Inf}}$ is the inferred causal strengths by IEE, GC, TE, or CCM. The PCC linearity of IEE ($0.724\pm 0.116$) is significantly higher than that of GC ($0.161\pm 0.099$), TE($0.284\pm 0.168$), and CCM ($0.568\pm 0.174$). (D) shows the GRN surrounding Clock, a key circadian gene. (E) demonstrates the ROC curves and AUC values of the IntDC index IEE, and three ObsDC indices GC/TE/CCM.
Application of IEE to investigating circadian rhythms
@@
We investigated the gene regulatory networks (GRNs) related to key genes on the circadian rhythm. The time series of gene expressions were measured by microarray from cultured rat cells  (?, ?). Through decades of molecular and genetic studies, many key circadian genes, such as Clock, Bmal1(Arntl), Dec1(Bhlhb2), Dec2(Bhlhb3), Cry1, Cry2, Per1, Per2, Per3, have been identified on mammals  (?, ?). Figure 6(D) displayed the GRN around Clock at the protein level, where the transcription factor Clock is phosphorylated by PFK family genes. The ROC curves and AUC values for IEE/GC/TE/CCM were showed in Fig. 6(E). IEE designed for quantifying IntDC presented a higher AUC value ($0.737$) than ObsDC indices ($0.490$ for GC, $0.610$ for TE, and $0.644$ for CCM). In fig. S14, we also tested the indices on the GRN surrounding Cry1/Cry2, where IEE outperformed the others.
Concluding remarks and perspective
@@
In summary, we have established a framework of IntDC based on dynamical systems theory and introduced the IEE criterion to quantify IntDC in this study. IEE measures the information flow between interventional causes and effects in the delay-embedding space, making it suitable for analyzing non-linear and non-separable systems (?, ?). Moreover, IEE is able to infer IntDC solely from observational time-series data, without requiring additional perturbations to the system. This property makes IEE particularly suitable for non-intervention systems. Both theoretical derivations and numerical examples presented in this study provide strong evidence to support the high effectiveness, accuracy and robustness of IEE in detecting IntDC and reconstructing networks. Furthermore, IEE serves as an effective tool for evaluating and ranking the causal dependence between variables within a dynamical system. Through real-world examples, we illustrated the promising applicability of IEE in diverse fields such as regulatory inference and disease transmission studies.
@@
There are several compelling issues that warrant further investigation and future development.
@@
(1) The theory of IntDC and the IEE criterion can be extended to analyze causal relations between groups of variables. Specifically, in Eq. (6), the variables $x$ and $y$ can be generalized to high-dimensional vectors, enabling the investigation of complex interactions among multiple variables.
@@
(2) The data in computing IEE can be non-uniformly sampled in the temporal space. We used the time-delayed embedding Eq. (2) and Eq. (3) in this paper, while according to the general Takens’ embedding theorem, any $L+1$ observations of $x_{t}$ and $y_{t}$ can form a topological diffeomorphism to $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$, respectively. Equation (6) shows that only neighbors in the embedding space matter. Thus, IEE algorithm can be extended to infer causality using diverse datasets sampled at different intervals.
@@
(3) To distinguish direct and indirect causality, we can generalize IEE to its conditional version
$\text{cIEE}[x\rightarrow y|z]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1},\bm{Z}_{t},\delta\bm{Z}_{t}),$ (10)
@@
where “cIEE” is short for the conditional IEE, $z$ is a third variable whose time-delay vector is $\bm{Z}_{t}$, and $\delta\bm{Z}_{t}$ represents the intervened deviation on $\bm{Z}_{t}$. Equation (10) measures the direct IntDC from $x$ to $y$ conditioned on $z$. Further, together with the well-known PC algorithm  (?), cIEE is one feasible way to remove high-order indirect causal edges iteratively and get the direct causal network.
(4) Moreover, causal inference with deep learning technology represents a frontier in research. Notable works include CausalEGM  (?) and Causalformer (?). Combining IntDC and neural networks is a promising field. Causality theory brings interpretability for artificial intelligence, while big model can provide new discoveries in causal relationships. We are now entering a brilliant “big causality era”.
@@ Acknowledgments
This work was supported by the National Key R&D Program of China 2022YFC2704604, the National Natural Science Foundation of China No. 12301620 and No. 32100509, the Moonshot R&D Grant Number JPMJMS2021, Institute of AI and Beyond of UTokyo, the International Research Center for Neurointelligence (WPI-IRCN) at The University of Tokyo Institutes for Advanced Study (UTIAS), JSPS KAKENHI Grant Number JP20H05921, the Cross-ministerial Strategic Innovation Promotion Program (SIP), and the 3rd period of SIP “Smart energy management system” Grant Number JPJ012207.
Author contributions:
Conceptualization: J.S., K.A., L.C., and W.L.
Methodology: J.S., K.A., L.C., and W.L.
Investigation: J.S., Y.L., J.Z., and S.L.
Visualization: J.S., Y.L., J.Z., S.L.
Supervision: K.A., L.C., and W.L.
Writing – original draft: J.S, Y.L., J.Z., and L.C.
Writing – review & editing: All authors.

### v2

@@ 3.6 Application of IEE to COVID-19 transmission in Japan
We validated that IEE was a promising indicator for assessing the transmission dynamics of infectious diseases. Our study collected the daily confirmed COVID-19 cases from all 474747 prefectures in Japan, spanning a duration of 1209 days from January 16, 2020 to May 8, 2023 (fig. S11).
@@
First, we calculated the IEE indices for IntDC among different prefectures. IEE effectively ranked the influence across prefectures, with the top five affected areas by Tokyo being Kanagawa (with IEE value 0.6010.6010.601), Chiba (with IEE value 0.5680.5680.568), Saitama (with IEE value 0.5340.5340.534), Aichi (with IEE value 0.4530.4530.453), and Osaka (with IEE value 0.4350.4350.435). Remarkably, this ranking coincides with the geographical proximity and socio-economic connections to Tokyo. Kanagawa, Chiba and Saitama are in the same metropolitan area with Tokyo, while Aichi and Osaka are far from Tokyo but connected to Tokyo by expressway and Shinkansen.
@@
Then, IEE was validated to coincide with the effective distance matrix $D^{\text{COVID}}$ between prefectures. The $D^{\text{COVID}}\in\mathbb{R}^{47\times 47}$ was non-symmetric and designed by incorporating factors such as geometric distances, human mobility, population sizes, and infectious rates across prefectures (based on the gravity model  (?, ?), see Supplementary Text and fig. S11 for details). The element $D_{ij}^{\text{COVID}}$ served as a benchmark for quantifying COVID-19 transmission from prefecture $i$ to $j$. The IEE values demonstrated a strong linear correlation with $\ln D^{\text{COVID}}$, particularly when assessing the causality from Tokyo/Osaka to other prefectures (with Pearson correlation coefficients, i.e. PCCs, 0.906-0.906- 0.906 for Tokyo and 0.860-0.860- 0.860 for Osaka, Figs. 6(A-B)). This correlation remained high for other regions (see fig. S12), such as Aichi (with PCC value 0.811-0.811- 0.811), Hokkaido (with PCC value 0.751-0.751- 0.751), Fukuoka (with PCC value 0.752-0.752- 0.752), and Okinawa (with PCC value 0.811-0.811- 0.811).
@@
Further, IEE outperformed ConDC indices in practicality. Notably, the absolute value of the PCC between IEE and $\ln D^{\text{COVID}}$ was 0.7240.7240.724 substantially higher than those obtained by GC 0.1610.1610.161, TE 0.2840.2840.284, and CCM 0.5680.5680.568 (average over specific prefecture, see the second to last column in Table 3 and the boxplot in Fig. 6(C)). More detailed comparisons to support the effectiveness of IEE were presented in Table 3 (with specific illustration for Tokyo in fig. S13).
@@
Figure 6: Application of the IEE on the COVID-19 transmission in Japan and the gene regulatory network (GRN) of circadian rhythms. (A) shows that IEE values from Tokyo to other prefectures are highly linear-correlated with $\ln D^{\text{COVID}}$, where $D^{\text{COVID}}$ is the effective distance obtained from the gravity model in consideration of various factors. (B) is the result for IEE values from Osaka to the other prefectures. (C) is the boxplot for $|\text{PCC}(c_{ij}^{\text{Inf}},\ln D_{ij}^{\text{COVID}})|$ with average over fixed $i$, where PCC is the Pearson correlation coefficient and $c^{\text{Inf}}$ is the inferred causal strengths by IEE, GC, TE, or CCM. The PCC linearity of IEE (0.724±0.1160.724$\pm$ 0.1160.724 ± 0.116) is significantly higher than that of GC (0.161±0.0990.161$\pm$ 0.0990.161 ± 0.099), TE(0.284±0.1680.284$\pm$ 0.1680.284 ± 0.168), and CCM (0.568±0.1740.568$\pm$ 0.1740.568 ± 0.174). (D) shows the GRN surrounding Clock, a key circadian gene. (E) demonstrates the ROC curves and AUC values of the IntDC index IEE, and three ConDC indices GC/TE/CCM.
3.7 Application of IEE to investigating circadian rhythms
@@
We investigated the gene regulatory networks (GRNs) related to key genes on the circadian rhythm. The time series of gene expressions were measured by microarray from cultured rat cells  (?, ?). Through decades of molecular and genetic studies, many key circadian genes, such as Clock, Bmal1(Arntl), Dec1(Bhlhb2), Dec2(Bhlhb3), Cry1, Cry2, Per1, Per2, Per3, have been identified on mammals  (?, ?). Figure 6(D) displayed the GRN around Clock at the protein level, where the transcription factor Clock is phosphorylated by PFK family genes. We calculated the IEE/GC/TE/CCM to reconstruct the GRN (ROC curves and AUC values in Fig. 6(E)). IEE designed for quantifying IntDC presented a higher AUC value (0.7370.7370.737) than ConDC indices (0.4900.4900.490 for GC, 0.6100.6100.610 for TE, and 0.6440.6440.644 for CCM). We also tested the indices on the GRN surrounding Cry1/Cry2 (fig. S14), where IEE obtained the highest AUC value (0.6390.6390.639) and outperformed the others (0.4280.4280.428 for GC, 0.6190.6190.619 for TE, and 0.6140.6140.614 for CCM).
4 Conclusions and discussions
@@
In summary, we have established a framework of IntDC based on dynamical systems theory and introduced the IEE criterion to quantify IntDC in this study. IEE measures the information flow between interventional causes and effects in the delay-embedding space, making it suitable for analyzing non-linear and non-separable systems (?, ?, ?). Moreover, IEE is able to infer IntDC solely from observational time-series data, without requiring additional perturbations to the system. This property makes IEE particularly suitable for non-intervention systems. Both theoretical derivations and numerical examples presented in this study provide strong evidence to support the effectiveness, accuracy and robustness of IEE in detecting IntDC and reconstructing networks. Furthermore, IEE serves as an effective tool for evaluating and ranking the causal dependence between variables within a dynamical system. Through real-world examples, we illustrated the promising applicability of IEE in diverse fields such as regulatory inference and disease transmission studies.
Information-flow-based methods have been developing rapidly in recent years  (?, ?, ?). Liang’s information flow theory, albeit with full nonlinearity in its theoretical formalism, is yet to be implemented for practical applications on real-world data. Existing implementations still rely on assumptions such as independent white noise and linearity when evaluating information flow from data. Our IEE algorithm, inspired by the conception information flow, is specifically designed to handle fully nonlinear systems and has demonstrated applicability across a variety of real datasets. The proposed IntDC framework, together with the IEE algorithm, advances the frontier of causality research by providing a practical and effective tool for nonlinear causal inference.
Runge et al.  (?) summarized recent advances and key challenges in the field of causal inference and developed the platform https://causeme.net, which serves as a valuable resource for researchers. Causal inference holds great promise for applications in earth system science and beyond. Our proposed IEE algorithm, together with the concept of IntDC, provides a novel and practical approach for investigating causality in complex nonlinear systems.
@@
There are several important open issues that warrant further investigation and future development.
@@
(1) The theory of IntDC and the IEE criterion can be extended to analyze causal relations between groups of variables. Specifically, in Eq. (2), the variables $x$ and $y$ can be generalized to multivariate vectors, enabling the investigation of complex interactions among multiple variables.
@@
(2) The data used for computing IEE can be non-uniformly sampled in time. In this study, we employed the standard time-delayed embedding Eq. (3) and Eq. (4). However, according to the generalized Takens’ embedding theorem, any set of $L+1$ observations of $x_{t}$ and $y_{t}$ can construct topological diffeomorphisms to the manifolds $\mathcal{M}_{X}$ and $\mathcal{M}_{Y}$, respectively. As shown in Eq. (7), only the local neighborhood structure in the embedding space is essential for the computation. Therefore, the IEE algorithm can be naturally extended to infer causality from datasets sampled at varying or irregular time intervals.
@@
(3) To further distinguish direct and indirect causality, we can generalize IEE to its conditional version
$\text{cIEE}[x\rightarrow y|z]:=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1},\bm{Z}_{t},\delta\bm{Z}_{t}),$ (11)
@@
where “cIEE” is short for the conditional IEE, $z$ is a third variable whose time-delayed vector is $\bm{Z}_{t}$, and $\delta\bm{Z}_{t}$ represents the intervened deviation on $\bm{Z}_{t}$. Equation (11) measures the direct IntDC from $x$ to $y$ conditioned on $z$. Further, in combination with the well-known PC algorithm  (?) or PCMCI (?), cIEE offers a feasible way to remove high-order indirect causal edges iteratively and reconstruct the direct causal network.
(4) In  (?, ?, ?), researchers have pointed out that time-delayed embedding algorithms are particularly suitable for deterministic dynamical systems with an attractive manifold. However, the coupling strength between variables could influence the accuracy of these methods. Stark et al.  (?) extended Takens’ embedding to stochastic systems, providing a theoretical basis for applying embedding approaches beyond purely deterministic dynamics. In the Results section of this study, we demonstrated the effectiveness of IEE for Logistic systems with Gaussian noises. Nevertheless, how systematic factors, such as coupling strength and noise characteristics, affect the quantification of causal strength in more general settings remains an open question.
(5) Moreover, causal inference empowered by deep learning is emerging as a frontier in research. Notable advances include CausalEGM  (?), Causalformer (?), intervened reservoir computing (?), and reservoir cross mapping (?). Integrating the IntDC framework with neural networks presents a promising direction for further exploration. On one hand, causality theory provides interpretability to artificial intelligence; on the other hand, large AI foundation models offer unprecedented capacity to uncover causal relationships in complex data . These developments signal the advent of a transformative “big causality era”, in which data-driven discovery and causal reasoning evolve in synergy.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China [grant numbers 42450192, 42450135, 12301620, 11925103 and 32100509], the National Key R&D Program of China [grant number 2022YFC2704604], the Shanghai Municipal Data Bureau special fund for urban digital transformation [grant number 202401065], the AI for Science Foundation of Fudan University [grant number FudanX24AI041], the AI for Science Program of Shanghai Municipal Education Commission [grant number 24RGZN10], the Japan Science and Technology Agency Moonshot R&D [grant number JPMJMS2021], the Institute of AI and Beyond of UTokyo, the International Research Center for Neurointelligence (WPI-IRCN) at The University of Tokyo Institutes for Advanced Study (UTIAS), the Cross-ministerial Strategic Innovation Promotion Program (SIP), the 3rd period of SIP [grant number JPJ012207].